A spot car skids off of a wet mountain at 48.5 m/s and lands in the river, 110 m below the road. How long will it take for the sports car to land in the water
1
Expert's answer
2020-12-21T10:43:47-0500

We can find the time that the car takes to reach the water from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 110\ m}{9.8\ \dfrac{m}{s^2}}}=4.74\ s."

Answer:

"t=4.74\ s."